eviant:

If f(x) = x2 -1, what is the equation for f^–1(x)?

eviant:

@dude

dude:

To solve for its inverse, you have to flip the x and y and then solve for y \(f(x) = x^2 -1\) f(x) is the same as y \(y = x^2 -1\) Now flip \(x=y^2-1\)

dude:

And then solve for y

eviant:

@dude how do I solve for y?

dude:

Add 1 to both sides

eviant:

x+1=y^2?

dude:

Yes Now do the square root on both sides

eviant:

x=-1+y^4?

dude:

wut

dude:

No \(\sqrt{x+1}=\sqrt{y^2}\)

eviant:

x=-1+y^2

dude:

I am not sure what you did It is just \(\sqrt{x+1}=y\) or \(y=\sqrt{x+1}\) The inverse is a square root function